Find the entries xi;i D 1; 2; 3; 4 in the matrix P so that P is an orthogonal matrix where
P D
1
2
2
6
6
4
1 1 1 1
1 1
1
Expert's answer
2020-04-07T13:01:16-0400

Let matrix "P" be


"P=k\begin{pmatrix}
 1 & a & b & c \\
 d & e & x_1 & x_2 \\
 1 & 1 & x_3 & x_4 \\
 1 & 1 & 1 & 1
\end{pmatrix}"

Let


"r_1=\begin{pmatrix}
 1 \\
 d \\
 1 \\
 1
\end{pmatrix},r_2=\begin{pmatrix}
 a \\
 e \\
 1 \\
 1
\end{pmatrix},r_3=\begin{pmatrix}
 b \\
 x_1 \\
 x_3 \\
 1
\end{pmatrix}, r_4=\begin{pmatrix}
 c \\
 x_2 \\
 x_4 \\
 1
\end{pmatrix}"

 "P" is an orthogonal matrix


"PP^T=I"

Then


"r_1r_2=r_1r_3=r_1r_4=r_2r_3=r_2r_4=r_3r_4=0""a+ed+1+1=0""b+dx_1+x_3+1=0""c+dx_2+x_4+1=0""ab+ex_1+x_3+1=0""ac+ex_2+x_4+1=0""bc+x_1x_2+x_3x_4+1=0"

Let "d=1." Then

"a+e+2=0""b+x_1+x_3+1=0""c+x_2+x_4+1=0""ab+ex_1+x_3+1=0""ac+ex_2+x_4+1=0""bc+x_1x_2+x_3x_4+1=0"

Solve


"a=-1, b=-1,c=1, d=1, e=-1,"

"x_1=1, x_2=-1,x_3=-1, x_4=-1"


"k^2r_1^2=k^2r_2^2=k^2r_3^2=k^2r_4^2=1"

"k=\pm0.5"


"P=0.5\begin{pmatrix}
 1 & -1 & -1 & 1 \\
 1 & -1 & 1 & -1 \\
 1 & 1 & -1 & -1 \\
 1 & 1 & 1 & 1
\end{pmatrix},P^T=0.5\begin{pmatrix}
 1 & 1 & 1 & 1 \\
 -1 & -1 & 1 & 1 \\
 -1 & 1 & -1 & 1 \\
 1 & 1 & -1 & 1
\end{pmatrix}"

"PP^T=0.5\begin{pmatrix}
 1 & -1 & -1 & 1 \\
 1 & -1 & 1 & -1 \\
 1 & 1 & -1 & -1 \\
 1 & 1 & 1 & 1
\end{pmatrix}\begin{pmatrix}
 1 & 1 & 1 & 1 \\
 -1 & -1 & 1 & 1 \\
 -1 & 1 & -1 & 1 \\
 1 & -1 & -1 & 1
\end{pmatrix}="

"=\begin{pmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1
\end{pmatrix}=I"

"x_1=1, x_2=-1,x_3=-1, x_4=-1"